OpenStudy (anonymous):

Solve for x ; X^-3=1/729 please help

OpenStudy (unklerhaukus):

\[x^{-3}=\frac1{729}\\[2ex] \frac1{x^3}=\frac1{729}\\[2ex]\quad x^3=729\\[2ex] \quad \ x=\sqrt[3]{792}\\[2ex]\qquad=\quad\dots\]

OpenStudy (anonymous):

@ig

OpenStudy (anonymous):

@iGreen